OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. A. y + 3= -6(x + 2) B. y + 2= 6(x + 3) C. y - 3= -6(x - 2) D. y - 2= 6(x - 3) Please help.

OpenStudy (anonymous):

ello love <3 :D haha need some help?

OpenStudy (anonymous):

@rebeccca_sonnenschein

OpenStudy (anonymous):

Yes/: please haha @atreyu6s3x6

OpenStudy (anonymous):

The basic point slope form is\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so the equation is y – y1 = m(x – x1) (x1,y1) is our point, and m is our slope :)

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

That is the formula and then you input your information into it. So you would get y-(-3)=-6(x-(-2)) and then simplify.

OpenStudy (anonymous):

okay so our point or the pair, is (x,y) so we plus that in for x1 and y1 and the m is our slope

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

bananaroof gave you the answer

OpenStudy (anonymous):

good job haha @Rebecca @bananaroof01